1)In a standard normal distribution, what proportion of the data set is within 1.75 standard deviations away from the mean? Write your answer in percent and with 2 decimal places. (Answer: 91.99 (with margin: 0.01) please show your work.)

2)A 100-point test was administered to a statistics class. Jack's score was in the 80th percentile. Suppose the statistics class has 120 students and all of them took the test, how many students scored lower than or equal to Jack's score?(Answer:96.0 (with margin: 0.0 please show your work.)

3) 12 18 22 23 24 25 26 27 29 30 32 33

Determine the Upper fence.(Answer: 40.0 (with margin: 0.0) and 41.0 (with margin: 0.0) please show your work.)
Answer #1

Solution:

Question 1) :

X ~ Standard Normal distribution with mean = μ=0 and standard deviation = \sigma = 1

We have to find proportion of the data set is within 1.75 standard deviations away from the mean.

That is: we have to find:

Pl-1.75 < X-μ < 1.75) = ?

1.75 X-μ 1.75 P(-1.75 < X-μ < 1.75)-P(

1.75 1.75 P(-1.75 < X-μ < 1.75)-P(

P( -1.75 < X - \mu < 1.75) = P( -1.75 <Z< 1.75 )

P( -1.75 < X - \mu < 1.75) = P( Z< 1.75 ) - P( Z < -1.75 )

Look in z table for z = 1.7 and 0.05 as well as for z= -1.7 and 0.05 and find area.

01 02 03 .04 05 06 .07 09 3.4 .0003 .0003 .0003 .0003 .0003 .0003 .0003 .0003 .0003 .0002 -3.3 .0005 .0005 .0005 .0004 .0004

P( Z < -1.75) = 0.0401

200 .010203.04 .05 .06 .07 08.09 0.0 .5000 5040 5080 .5120 .5160 .519 0.1 5398 .5438 .5478 .5517 .5557 .5596 .5636 .5675 .571

P( Z< 1.75) = 0.9599

Thus

P( -1.75 < X - \mu < 1.75) = P( Z< 1.75 ) - P( Z < -1.75 )

P(-1.75 < X-μ < 1.75) = 0.9599-0.0401

P( -1.75 < X - \mu < 1.75) =0.9198

To get exact answer as 0.9199, we need to use Calculator or Excel command:

Excel command is:

=NORM.S.DIST( 1.75 , TRUE) - NORM.S.DIST( - 1.75 , TRUE)

= 0.9199